OpenStudy (anonymous):

*HELLLLPPP* A colony of bacteria is growing exponentially according to the function below, where t is in hours. How many hours will it take for the population to grow to 500? Round to the nearest integer. B(t) = 10* e^t

OpenStudy (anonymous):

B(t)=500 Plug in to the equation: 500=10e^t e^t = 50 lne^t = ln (50), which cancels out the e^ because lne = 1 t= ln( 50), plug in calculator to find answer 3.912 hr round to nearest 100th decimal places.

OpenStudy (anonymous):

thank you so much! so, the answer would be 4 hours?

OpenStudy (anonymous):

Yes ,if rounded to nearest whole number c:

OpenStudy (anonymous):

Thank you!!!! :)